# Shared DIFF Transformer

### v1

@@ 3.5 Ablation Studies
We conduct ablation studies using the same training setup as described for the 1.4B model in Section 3.2. Table 6 reports the fine-grained loss on the validation set, breaking it into two components: ”AR-Hit” and ”Others.” ”AR-Hit” evaluates the model’s ability to recall previously seen n-grams, while ”Others” represents tokens that are either frequent or not recalled from the context.
As shown in Table 6, we conducted ablation studies on various design choices in DIFF Transformer and Shared DIFF Transformer and compared their performance. The first and fifth rows correspond to the default settings for DIFF Transformer and Shared DIFF Transformer, respectively. The results clearly show that GroupNorm has a significant impact on the model’s performance, as the rows in the attention matrices of both models no longer sum to 1. GroupNorm ensures numerical stability in this case.
We also experimented with different $\lambda$ initialization strategies. The results indicate that the default initialization method outperforms the constant initializations of $\lambda_{\text{init}}=0.8$ and 0.5. Moreover, the performance gap is minimal, highlighting the effectiveness of the initialization strategy and the robustness of the model to different initialization methods.

### v2

@@ 3.5 Ablation Studies
We conducted ablation studies with the same training setup as in Section 3.2 for the 1.4B model. Table 4 reports the fine-grained loss on the validation set, divided into ”AR-Hit” (recall of n-grams) and ”Others” (frequent or unrecalled tokens).
As shown in Table 4, ablation studies on various design choices in DIFF and Shared DIFF Transformers reveal that GroupNorm significantly impacts performance, ensuring numerical stability by making the attention matrices sum to 1. We also tested different $\lambda$ initialization strategies, finding that the default method outperforms $\lambda_{\text{init}}=0.8$ and 0.5, with minimal performance differences, demonstrating the robustness of the model to initialization variations.